OpenStudy (anonymous):

The preimage of C'(-3,-10) use the translation (x,y)(x-8,y+4)

jimthompson5910 (jim_thompson5910):

They want to know a point, call it C, where you apply the translation rule to point C to get C'

jimthompson5910 (jim_thompson5910):

So you're going from C(x,y) to C'(-3,-10)

jimthompson5910 (jim_thompson5910):

using the rule (x,y) --> (x-8, y+4)

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Do you see what to do to find C(x,y) ?

OpenStudy (anonymous):

Plug in (-3,-10) ?

jimthompson5910 (jim_thompson5910):

into what?

OpenStudy (anonymous):

(x-8,y+4) ?

jimthompson5910 (jim_thompson5910):

that will apply the rule to C' to get C'' but we want to undo that rule to go back to C

jimthompson5910 (jim_thompson5910):

you just need to solve these two equations x-3 = -8 y+4 = -10 the (x,y) pair you get will be the point C

OpenStudy (anonymous):

I keep getting -5,-14 but the answer is 5,-14

jimthompson5910 (jim_thompson5910):

x-3 = -8 x-3+3 = -8+3 x+0 = -5 x = -5

jimthompson5910 (jim_thompson5910):

So it should be (-5,-14)

jimthompson5910 (jim_thompson5910):

oh wait

jimthompson5910 (jim_thompson5910):

nvm

jimthompson5910 (jim_thompson5910):

I wrote down the equation wrong

jimthompson5910 (jim_thompson5910):

it should be x-8 = -3 x-8+8 = -3+8 x+0 = 5 x = 5

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

you're welcome